### v1

# 2.5D Transformer: An Efficient 3D Seismic Interpolation Method without Full 3D Training

Abstract
Transformer has emerged as a powerful deep-learning technique for two-dimensional (2D) seismic data interpolation, owing to its global modeling ability. However, its core operation introduces heavy computational burden due to the quadratic complexity, hindering its further application to higher-dimensional data. To achieve Transformer-based three-dimensional (3D) seismic interpolation, we propose a 2.5-dimensional Transformer network (T-2.5D) that adopts a cross-dimensional transfer learning (TL) strategy, so as to adapt the 2D Transformer encoders to 3D seismic data. The proposed T-2.5D is mainly composed of 2D Transformer encoders and 3D seismic dimension adapters (SDAs). Each 3D SDA is placed before a Transformer encoder to learn spatial correlation information across seismic lines. The proposed cross-dimensional TL strategy comprises two stages: 2D pre-training and 3D fine-tuning. In the first stage, we optimize the 2D Transformer encoders using a large amount of 2D data patches. In the second stage, we freeze the 2D Transformer encoders and fine-tune the 3D SDAs using limited 3D data volumes. Extensive experiments on multiple datasets are conducted to assess the effectiveness and efficiency of T-2.5D. Experimental results demonstrate that the proposed method achieves comparable performance to that of full 3D Transformer at a significantly low cost.
keywords:
Seismic data interpolation, Cross-dimensional transfer learning, Transformer, Seismic dimension adapter1 Introduction
Interpolation operation is a crucial step for improving the quality of seismic data, especially when facing geometries with large spatial sampling intervals, and recovering missing parts caused by topographical constraints, limited budgets, or receiver malfunctions. By reconstructing the missing information and compensating the sparse geometries, effective interpolations can enhance the completeness and continuity of events, facilitating the subsequent inversion and high-resolution imaging [chen2024low, cheng2025generative].
To acquire the desired dense and complete seismic data, geophysical scholars have proposed various conventional interpolation methods over the past few decades [spitz1991seismic, trad2002accurate, oropeza2011simultaneous]. These methods are theory-driven, relying on mathematical models or analytical formulations to describe the underlying physical processes. Generally, they can be divided into four categories: the wave-equation-based, prediction filtering, low-rank, and sparse-representation methods.
The wave-equation-based methods utilize subsurface velocity models to simulate the propagation of seismic wavefields, so as to reconstruct the incomplete data [wang2019deep]. [ronen1987wave] proposed a trace-interpolation method based on the wave equation and prior assumption of a smooth spatial spectrum, and experimental results have demonstrated its effectiveness on both synthetic and field data. The offset-continuation differential equation-based method proposed by [fomel2003seismic] showed good interpolation performance even in several structurally complex situations. However, the wave-equation based methods often require a relatively accurate velocity model as an important prerequisite, which is often unavailable in numerous real-world problems. Moreover, the heavy computation burden seriously limits their further applications.
The prediction filtering methods exploit the correlation of events in different domains to perform the interpolation operation. When applying these methods, the interpolation task is taken as a least-squares linear inverse problem, and interpolated results are generated by minimizing the misfit between the predicted and input data [liu2022bseismic]. [spitz1991seismic] proposed a multichannel and model-free interpolation method for missing traces in f-x domain, exhibiting good interpolation performance in both pre- and post-stack examples. To avoid solving the complex linear system of equations in Spitz method, [porsani1999seismic] introduced a half-step predictive filter to make the interpolation operation more efficient and easier for implementation. [wang2002seismic] extended the f-x domain interpolation to f-x-y domain by masking the data samples selectively from the input frequency slice to design the linear prediction (LP) operators. To further enhance the interpolation performance of predictive filters on irregularly sampled data, [naghizadeh2007multistep] proposed a multistep autoregressive algorithm-based prediction-error method combined with Fourier-based methods to reconstruct seismic data from low to all frequencies. An important assumption of these prediction filtering methods is that events are required to be linear. However, this is usually not satisfied in real conditions, resulting in serious interpolation performance degradation.
The low-rank methods assume that the complete seismic data in a specific arrangement is of low rank at a given frequency component [trickett2010, gao2013fast], and missing traces will significantly increase its rank. Therefore, the interpolation tasks can be accomplished by reducing the rank to an optimal one. Typical low-rank methods include singular spectrum analysis [SSA; vautard1989singular, oropeza2011simultaneous, lari2019adaptive, carozzi2021interpolated, li2024interpolated], matrix completion [ma2013three, yang2013seismic, kumar2015efficient], principal component analysis [WOLD198737, li2015inverse, huang2016improved, wu2023high], and Cadzow filtering [cadzow2002signal, gao2013fast, naghizadeh2013multidimensional, huang2020dealiased]. [oropeza2011simultaneous] proposed a simultaneous denoising and reconstruction method for seismic data based on multicahnnel SSA (MSSA), which resembles seismic data interpolation with the method of projection onto convex sets (POCS). Moreover, a randomized singular value decomposition is adopted to accelerate its rank reduction stage. [ma2013three] enhanced the matrix completion using a designed texture-patch transformation, exhibiting superior performance to traditional POCS method. [naghizadeh2013multidimensional] used the low-rank approximation of the Hankel matrix of seismic data to recover missing spatial samples. Its effectiveness was validated in the reconstruction of synthetic and real data examples. However, in these low-rank methods, the determination of optimal rank is an intricate problem, and an improper one will have a negative impact on the interpolated results [ma2013three, cheng2023seismic].
The sparse-representation methods interpolate seismic records by representing them into sparse domains. The useful signals and the missing data are treated as large- and small-amplitude coefficients, respectively. Preserving the large-amplitude coefficients and eliminating the small-amplitude ones helps to extract the useful signal components from the incomplete data [chen2019interpolation]. The missing signal components are then recovered through an inverse sparse transform. These methods can be mainly categorized into mathematical transform-based and dictionary learning-based methods. [trad2002accurate] proposed a high-resolution time-variant Radon transform (RT)-based interpolation method. Hyperbolic and elliptical RTs are implemented to perform accurate interpolation and attenuate sampling artifacts in poorly sampled common-midpoint gathers. [yu2007wavelet] proposed a wavelet-Radon transform-based dealiasing algorithm that exploited the blurring of frequency information across wavelet scales to interpolate aliased signal components, exhibiting better performance than the Spitz method does. A new digital transform named seislet transform was proposed by [fomel2010seislet] to interpolate seismic data. It combined the wavelet-lifting scheme with local plane-wave destruction, providing a convenient orthogonal basis for denoising and interpolation. [naghizadeh2010beyond] proposed an interpolation method based on the curvelet transform for aliased seismic data. This study proves that spatially aliased data can be represented in the curvelet domain by two types of coefficients in coarser and finer scales. As a newly developed sparse transform at the time, the dreamlet transform was adopted by [wang2014dreamlet] to interpolate synthetic and field seismic data, exhibiting better performances in terms of sparsity and signal-to-noise ratio (SNR) than those of the curvelet transform-based method. Based on the good sparse representation ability of shearlet transform, [liu2018crossline] proposed a multi-component crossline seismic data reconstruction method based on sparse shearlet constraint inversion, achieving better results than traditional wavelet, curvelet, and shearlet methods in several extremely sparse sampling cases. Most of the above mathematical transform-based methods are based on fixed bases [liang2014seismic]. To learn basis functions adaptively, experts have developed dictionary learning-based interpolation methods. [liang2014seismic] attempted to restore decimated seismic data using data-driven tight frame (DDTF) first developed by [cai2014data]. The DDTF can adaptively learn from the data itself, and provides a sparser representation for the data in turn. [yu2015interpolation] extended the DDTF to high-dimensional versions and achieves the simultaneous denoising and interpolation of 3D and 5D seismic data. [wang2020fast] proposed an efficient dictionary learning method for high-dimensional seismic data reconstruction based on sequential generalized K-means model instead of the K-singular value decomposition (KSVD), achieving comparable interpolation performance to the traditional KSVD method at a very low computational cost. Nevertheless, these sparse-representation methods are hampered by inherent limitations, despite their excellent anti-aliasing and interpolation performance. First, the sparsity of data is an essential assumption. Second, the choice of parameters plays a crucial role. For instance, for the curvelet transform, inaccuracies in the estimation of the mask function can directly affect interpolation performance; for the seislet transform, inaccurate estimation of the local event slope tends to deteriorate the fidelity of seismic data reconstruction [wang2019deep].
Generally, the successful applications of these conventional methods must satisfy a certain number of prior assumptions, including an accurate subsurface velocity model, the linearity of events, the low-rank structures of seismic data, and the sparsity of signals [jia2017what, wang2019deep, dong2025deep]. However, it is usually difficult to ensure that the understanding of the subsurface structure is consistent with real conditions [cheng2024meta], restricting the effectiveness and adaptability of these theory-driven interpolation approaches in handling several real and complex scenarios [jia2017what]. These theory-driven methods also involve complex and sophisticated operations for parameter adjustments, which extremely requires expert knowledge and hands-on experience. Additionally, the heavy computational cost that occasionally occurs is also a major hurdle, especially when handling with large-scale and high-dimensional seismic datasets [wang2019deep, dong2022seismic].
In recent years, data-driven techniques have attracted much attention for their ability to learn from the data itself and process large amounts of data more efficiently [cheng2024meta]. As a representative data-driven technique, deep learning [DL; lecun2015deep] refers to a class of representation-learning methods that automatically mine useful features from raw data by stacking numerous nonlinear transformation layers. Each layer extracts increasingly abstract representations, enabling the system to learn complex functions directly from the input data without manual feature engineering [lecun2015deep, alzubaidi2021review, mousavi2022deep]. These characteristics enable DL-based methods perform well without prior assumptions and handcrafted parameters, so as to automatically process large-scale datasets. Following the pioneering efforts [mandelli2018seismic, mikhailiuk2018deep, wang2018seismic], researchers have increasingly explored DL-based approaches for seismic data interpolation, with convolutional neural networks (CNNs) being the primary focus in the early phase. Typical CNN-based architectures, such as U-Net [ronneberger2015unet, park2019reconstruction, park2021method, fang2021seismic], Res-Net [He_2016_CVPR, wang2018seismic, wang2019deep, liu2022aseismic], and generative adversarial network [GAN; goodfellow2014generative, siahkoohi2018seismic, oliveira2018interpolating], have shown great promise for the seismic data interpolation. Various improvements applied to CNNs have also contributed to further enhance the interpolation performance, including the attention-enhanced CNNs [yu2021attention], depthwise separable CNNs [jin2023depthwise], and multi-scale CNNs [cheng2023seismic, dong2024can, dong2024seismic]. As researchers have gained deeper insights into DL-based interpolation methods, the local perception has emerged as a major bottleneck of CNNs. This property leads to the neglect of global features and hinders the further interpolation performance enhancement using DL. To address this issue, experts have turned to Transformer [vaswani2017attention] owing to its strong ability to extract global contextual information using self-attention mechanism. [guo2023seismic] proposed to apply the window-based multi-head self-attention (MSA) into a multi-scale Transformer network. It can capture long-range dependencies to reconstruct more detailed information than CNNs do. Utilizing shifted window based MSA, [peng2023irregularly] trained an interpolation Swin-Transformer that combines Swin-Transformer modules and dense skip connections in a self-supervised manner, achieving higher accuracy and robustness than the convolutional autoencoder method. [cheng2024seismic] proposed a seismic interpolation Transformer network, which is an encoder-decoder structure, with Swin-Transformer blocks forming a U-shaped structure between the encoder and decoder. This approach can effectively reconstruct the consecutively missing traces in distributed acoustic sensing-vertical seismic profiling data. By integrating the global and local features, [gao2024swin] propose a Swin-Transformer convolutional residual network for the simultaneous denoising and interpolation of seismic data. This CNN-Transformer hybrid model has gain better visual performance and quantitative indices than conventional methods and CNN-based methods. A dense double branch attention Transformer (D2AT) proposed by [dong2025deep] can effectively reconstruct the consecutively missing traces. The core module of D2AT is a global feature unit composed of six Swin-Transformer blocks distributed in two scales, and it has shown great interpolation performance and generalization in both pre-stack and post-stack datasets.
Although these DL-based methods have achieved superior performance over conventional methods, they primarily focus on two-dimensional (2D) interpolation and overlook the spatial structure correlation that could be exploited in a 3D interpolation workflow [wang2019deep, liu2022aseismic, dong2025deep]. However, 3D feature extraction operations usually introduce significantly greater computational burdens for DL-based interpolations. This phenomenon is particularly prominent when applying Transformer-based architectures, whose computational complexity scales quadratically with the size of input data [vyas2020fast]. Moreover, extending 2D networks to 3D versions will significantly increase the number of trainable parameters [ye20193d] . Several scholars have made attempts on 3D seismic data interpolation based on 3D CNNs in recent years [qian2021dtae, chen2023projection, saad2023unsupervised, wang2025self], but most of these methods suffer from heavy computational burdens. Obviously, this phenomenon will become more severe in 3D Transformers due to their quadratic computational complexity. To the best of our knowledge, no prior effort has attempted to apply Transformers to 3D seismic data interpolation at this stage, as training a 3D Transformer network on 3D seismic datasets demands prohibitively large computational resources. The application of Transformers to 3D seismic data interpolation remains an uncharted research area and holds significant potential for future investigations.
A recent study [pan2022st] demonstrates a promising approach to address the above challenges. [pan2022st] proposed an image-to-video transfer learning (TL) strategy that can transfer an 2D image model to a 3D video model using parameter-efficient fine-tuning operations. Inspired by this cross-dimensional (i.e., 2D to 3D) TL strategy, we propose a lightweight 2.5-dimensional (2.5D) Transformer (T-2.5D) network to reconstruct 3D incomplete seismic data without using a large amount of 3D volumes to optimize a 3D Transformer network. The proposed T-2.5D is a hybrid of 2D and 3D modules, including four 2D Transformer encoders and four 3D seismic dimension adapters (SDAs). We design a 2D-to-3D cross-dimensional training workflow to optimize the T-2.5D. This workflow contains two stages: a 2D pre-training stage and a 3D fine-tuning stage. In the 2D pre-training stage, we use 2D training patches to just optimize the four 2D Transformer encoders without SDAs. In the 3D fine-tuning stage, we freeze the trainable parameters of the four 2D Transformer encoders and use limited 3D volumes to train the four 3D SDAs. This fine-tuning stage enables the T-2.5D to learn spatial correlation information of 3D seismic data. Overall, we use the ‘2D pre-training + 3D fine-tuning’ to replace the time-consuming and memory-intensive full 3D training, so as to significantly alleviate the computational cost of 3D Transformer-based interpolation. We investigate the interpolation performance of T-2.5D on several 3D seismic volumes. Experimental results show that the T-2.5D achieves comparable interpolation performance to 3D Transformer at a significantly lower computational cost. In other words, the proposed T-2.5D achieves a better trade-off between interpolation performance and computational cost.
The remainder of this paper is organized as follows. We begin with the theory of Transformer. Then, detailed descriptions of the proposed T-2.5D and the cross-dimensional TL are provided. Subsequently, we present the experimental results of several field examples. Finally, we draw a conclusion by summarizing our work.
2 Methodologies
In this section, detailed descriptions of the Transformer, Transformers with different dimensions, cross-dimensional TL, and Transformer-based interpolation theory are provided in the following subsections.
2.1 Transformer
As a classical DL framework, Transformer has gained significant attention from academia and industry in recent years. It has been applied to the fields of natural language processing [vaswani2017attention] and computer vision [dosovitskiy2020image]. Owing to the core operation, MSA, Transformer is capable of capturing the global contextual information and shows better performance than CNNs in numerous cases.
To prepare a 1D input sequence for the Transformer encoder shown in Fig. 1a, we first flatten the 2D input matrix using positional encoding and shape transformation (i.e., the 2D Handler in Fig. 1a). Then, the flattened feature is input into a layer normalization (LN) layer and an MSA layer, and it is also added to the original input using a residual connection. The output feature of MSA is then fed into another LN layer and a multi-layer perceptron (MLP) layer, and is add to the final output feature via the second residual connection. The output of Transformer encoders is expressed as follows:
where and are the input and output features of Transformer encoders, and represents the 2D handler including the positional encoding and shape transformation. For brevity, or 3D handler is omitted in subsequent descriptions.
The implementation of MSA is illustrated in Fig. 1b. To compute the self-attention of the -th head, the input feature is linearly projected into queries (), keys () of channel number , and values () of channel number . First, we transpose the vector . Second, after the matrix multiplication between and , the product is divided by and then input into a softmax function to generate the weights for . Finally, the attention value is obtained by multiplying with its weights. The process is described by the following equation:
In order to fuse the information from various representation sub-spaces, the single-head attention is expanded to multi-head attention by projecting s, s and s for times using separate learnable linear layers. These attentions are then integrated using concatenation operation. Finally, a linear projection is utilized to project the channel number back to the same as the input. The MSA is calculated as follows:
2.2 Transformers with different dimensions
In this paper, we have designed three Transformer-based networks with different dimensions, including a full 2D Transformer (T-2D), a full 3D Transformer (T-3D) and a 2.5D Transformer (T-2.5D). They are highly similar in terms of architecture to ensure clear and relatively fair comparisons. Detailed descriptions of these three networks are provided below.
2.2.1 T-2D/3D
In Fig. 2a, the T-2D is composed of two 3×3 convolutional layers (Convs), a head block (HB), four 2D Transformer encoders, and a tail block (TB). A 2D patch with a size of 40×40 is firstly input into the first Conv, and the channel number is increased from 1 to 32. Then, the features with a size of 40×40×32 are input into HB whose structure is shown in Figure 2b. The HB consists of two identical parts, and each one comprises two Convs, a rectified linear unit (ReLU), and a residual connection. The output features of HB are propagated into the core module of T-2D: four successively-connected Transformer encoders, thereby capturing long-range dependencies. Moreover, dense connections [Huang_2017_CVPR], whose effectiveness has been validated by [dong2025deep], are deployed to enrich the feature interactions among the four Transformer encoders. Subsequently, the TB displayed in Fig. 2c is used for final feature refinement, which is composed of three Convs interleaved with leaky ReLUs in a feed-forward manner. The TB enables efficient non-linear transformation while mitigating the risk of neuron inactivation. The output features of TB and HB are fused via a residual connection, thus avoiding the phenomenon of gradient vanishing. Finally, to generate the final output feature, the last Conv is used to integrate features across channel dimension and reduce the channel number back to one.
We replace all replaceable layers of T-2D with their 3D counterparts to generate the T-3D shown in Fig. 3.
2.2.2 T-2.5D
The T-2.5D is generated by adding four 3D SDAs to T-2D. In Fig. 4a, we place a 3D SDA before each 2D Transformer encoder. As shown in Fig. 4b, the 3D SDA is composed of two linear layers and a depth-wise 3D Conv (DWConv3d). Specifically, the first linear layer is used to project the input feature into a 3D space, allowing the following DWConv3d to capture abundant spatial contextual information. At the end, another linear layer projects the feature back to the original dimension.
2.3 Cross-dimensional TL
We utilize a cross-dimensional TL strategy to optimize the T-2.5D. This workflow comprises two stages, including a 2D pre-training one and a 3D fine-tuning one. In the first stage, we optimize the four 2D Transformer encoders of the T-2.5D using a large number of 2D patches, so as to generate a 2D pre-trained model. In the second stage, we freeze the trainable parameters of the four Transformer encoders and use a small amount of 3D volumes to optimize the four 3D SDAs of T-2.5D, where spatial information across seismic lines is learned. The overall process of the cross-dimensional TL is provided in Fig. 5.
2.4 Transformer-based Interpolation Theory
2.4.1 2D/3D Interpolation
For the complete 2D seismic data , its corresponding incomplete 2D data can be expressed as:
where represents a 2D masking matrix used to generate the decimated data with missing traces. Through the network training of T-2D, we can establish a nonlinear mapping relationship between and , and the trainable parameters of T-2D is updated by minimizing the L2-norm loss function in Eq. (7).
where and denote the weights and biases, respectively, is the batch size (i.e., the pair number of training data patches used in one iteration), and represent the -th paired training data patches, and denotes the L2 norm. In this paper, we use the ADAM algorithm [kingma2014adam] to update the trainable parameters in an iterative manner. Upon obtaining the optimal parameters , we can generate the interpolated result by inputting the incomplete data into the nonlinear relationship :
The principle of 3D interpolation is similar to the 2D case, where we utilize 3D data to optimize the 3D trainable parameters of T-3D.
2.4.2 2.5D Interpolation
In this paper, we utilize two-stage training strategy to optimize the trainable parameters of T-2.5D, , so as to achieve the 2.5D interpolation using Transformer. The is composed of , , representing the trainable parameters of the 2D Transformer encoders and 3D SDAs, respectively.
In the 2D pre-training stage, the loss function is updated as follows:
where and denote the -th paired 2D training data patches, respectively. Upon obtaining the optimal parameters , we freeze them and use a small amount of 3D volumes to optimize the 3D SDAs. In the 3D fine-tuning stage, the loss function is updated through Eq. (10).
where and represent the i-th paired 3D training data volumes, respectively, and denotes the frozen parameters of 2D Transformer encoders. After certain iterations of the fine-tuning stage, we can obtain a set of optimal 2.5D parameters . The final 3D interpolation result is generated by inputting the incomplete 3D data into the 2.5D nonlinear relationship as follows:
The specific process of 2.5D interpolation is given in Algorithm 1.
3 Experiments
3.1 Training and Hyperparameters Settings
The training program is executed on Pytorch 2.5.1 and CUDA 12.4, running on an Ubuntu 22.04 operation system. Hardware configurations consist of an NVIDIA L20 GPU with 48GB of memory, an Intel Xeon Platinum 8457C CPU with 48 cores at 2.6GHz frequency, and 100GB RAM. Taking into account performance, computational resources, and fair comparison, we set the hyperparameters as shown in Table 1.
| Hyperparameters | Specifications |
| Optimizer | ADAM |
| Loss function | L2 norm |
| Data patch size | 40×40 (T-2D), 40×40×16 (T-3D and T-2.5D) |
| Batch size | 1 |
| Number of epochs | 20 |
| Learning rate | 10-4, 10-5 |
| Input channel number | 1 |
| Embedding channel number | 32 |
| Total layers | 4 |
[…]
3.2 Metrics for Interpolation Results
In this study, the peak signal-to-noise ratio (PSNR) and structural similarity index measure [SSIM; wang2004image] are used to evaluate the performance of different methods quantitatively. The PSNR of 2D data is defined as the following equation:
where and denote the complete and interpolated data, respectively, represents the number of sample points in each trace, is number of traces, and represents the maximum value of . Similarly, the 3D version of PSNR is expressed as follows:
where , , and denote the number of sampled points along the time, inline, and crossline axes, respectively.
SSIM is a commonly used metric to evaluate the perceptual similarity between two data. It considers the changes in structural information, luminance, and contrast [wang2004image]. Given a 2D/3D complete data and its corresponding interpolated result , SSIM is defined as:
where and represent the means of and , respectively, and are the variances of and , respectively, is the covariance between and , and , are small constants to stabilize the division of whole equation. The value of SSIM ranges from 0 to 1, with higher value indicating greater similarity.
3.3 Test of Kerry Dataset
3.3.1 Data preparation and training
We firstly use the Kerry dataset obtained from New Zealand to evaluate the effectiveness and efficiency of T-2.5D. The time sampling interval is 0.004 s. We extract a volume of 256 time samples, 650 inline points, and 186 crossline points from the Kerry dataset as shown in Fig. 6a. The training and test blocks are defined by red and green double arrows in Fig. 6a, respectively. In T-2D, we extract 7500 2D slices with a shape of 40×40 from the training block, and split them into training and validation sets in a ratio of 4:1. Similarly, in T-3D, 7500 3D volumes (40×40×16) are extracted from the training block and partitioned into training and validation sets with the same 4:1 ratio. In the first stage of T-2.5D, we pre-train the 2D Transformer encoders using the same dataset as in T-2D. The second stage of T-2.5D utilizes 2000 and 500 3D volumes as training and validation sets, respectively, to fine-tune the 3D SDAs. We randomly remove 40%–60% of the traces in each complete data patch/volume to generate complete-incomplete data pairs. Fig. 6b–6g displays the L2 loss curves for both training and validation of the three methods. After training, we input the test data into the three trained models to evaluate the interpolation performance. Notably, the interpolated result of T-2D is generated by processing the test data line by line and reassembling the outputs into a 3D volume.
3.3.2 Testing results
As shown in Fig. 7a, the test data is of shape 224×224×20 extracted from Fig. 6a. We randomly delete 50% of the traces to generate the incomplete data in Fig. 7e. Fig. 7b-7d presents the reconstruction results. Fig. 7f–7h display the residual images between the interpolated results (Fig. 7b–7d) and the complete data (Fig. 7a), and Fig. 7i–7k are the local similarity [fomel2007local] maps between Fig. 7f–7h and Fig. 7a. Both of them are utilized to measure the signal leakage of the interpolated results, and lower energy indicates better interpolation performance. We can observe that all the three methods are able to reconstruct the missing traces to some extent. However, T-2D exhibits stronger signal leakage than T-3D and T-2.5D as shown in Fig. 7f and 7i. In comparison, T-2.5D achieves a comparable performance to T-3D, which validates the effectiveness of T-2.5D.
For a clearer comparison, the sixth line is plotted in 2D (Fig. 8). Due to the lack of spatial correlation information, T-2D yields the poorest performance. This is evidenced by its significantly lower quantitative results and the most severe signal leakage among all networks, as shown in Fig. 8c and 8d. In comparison, T-2.5D achieves an interpolation result close to that of T-3D, and both of their residual images show visually little signal leakage.
Furthermore, we plot in Fig. 9 the f-k spectra of the complete data (Fig. 8a), 50% randomly sampled data (Fig. 8b), the three interpolated results (Fig. 8c, 8e, and 8g), and the corresponding residual images (Fig. 8d, 8f, and 8h). As illustrated in Fig. 9c–9e, the aliased energy presented in corrupted data (Fig. 9b) has been largely removed by all the three methods. However, as shown in Fig. 9c, stronger residual interference exists in the result of T-2D. It is also observed in Fig. 9h that the energy leakage of T-2.5D is evidently weaker than that of T-2D, and is very close to that of T-3D.
3.3.3 Analysis of the computational cost
Computational cost plays a crucial role in DL-based methods as it directly reflects the efficiency of different methods. We have recorded the peak memory usage and training time of the three networks in Table 2. Notably, the total training time of T-2.5D is composed of the pre-training time of the first stage and the training time of the second stage. T-2.5D requires only about 1/12 the memory of T-3D with the training time reduced to approximately 1/20. Generally, T-2.5D achieves comparable interpolation performance to T-3D, while substantially reducing the memory usage and training-time cost, demonstrating that T-2.5D is an efficient DL-based seismic data interpolation method and the proposed cross-dimensional TL is a lightweight process.
| Method | Peak memory usage (MB) | Training time (h) |
| T-2D | 581 | 0.46 |
| T-3D | 40443 | 19.90 |
| T-2.5D | 3243 | 0.46 (Stage 1) + 0.65 (Stage 2) |
[…]
3.4 Test on Parihaka Dataset
3.4.1 Data preparation and training
In this subsection, we use the Parihaka dataset obtained from another survey area of New Zealand to further validate the effectiveness of T-2.5D. As shown in Fig. 10a, the data preparation process of the Parihaka dataset is the same as that of the Kerry dataset. The data preparation of the three methods are given in Table 3. Fig. 10b–10g shows the convergence behavior of the training and validation curves of the three methods on the Parihaka dataset.
| Dataset size | |||
| Method | Data patch/volume shape | Training | Validation |
| T-2D | 40×40 | 6000 | 1500 |
| T-3D | 40×40×16 | 6000 | 1500 |
| T-2.5D | 40×40×16 | 2000 (Stage 2) | 500 (Stage 2) |
[…]
3.4.2 Testing results
The 3D data volume (224×224×20) defined by green double arrows in Fig. 10a is likewise extracted for testing. As shown in Fig. 11b–11d, the performance of T-2D remains markedly inferior to that of T-3D and T-2.5D, which is also supported by the strongest signal leakage in its residual image (Fig. 11f) and local similarity map (Fig. 11i). In contrast, the performance of T-2.5D is very close to that of T-3D, as shown in Fig. 11h and 11k.
Similarly, the tenth line is illustrated in Fig. 12 as a 2D view. The numerical results in Fig. 12g and little leakage in Fig. 12h indicate the great interpolation ability of T-2.5D.
Fig. 13 displays the f-k spectrum of the complete data (Fig. 12a), 50% randomly sampled data (Fig. 12b), the interpolated results in Fig. 12c, 12e and 12g, respectively, and the corresponding residual images (Fig. 12d, 12f, and 12h). T-2D still produces more residual interference than the other two methods as indicated by the red arrows and ovals in Fig. 13c. In comparison, the recovered spectrum in Fig. 13e and the minimal energy leakage in Fig. 13h demonstrate the strong interpolation capability of T-2.5D.
3.4.3 Analysis of the computational cost
Similar to Kerry dataset, we also provide comparisons of computational cost in Table 4. Due to the same settings of dataset preparation and training, the values on the Parihaka dataset are very similar to that of Kerry dataset. The T-2.5D has achieved great performance at very low computational cost, offering a good trade-off between the effectiveness and efficiency.
| Method | Peak memory usage (MB) | Training time (h) |
| T-2D | 581 | 0.44 |
| T-3D | 40443 | 19.94 |
| T-2.5D | 3243 | 0.44 (Stage 1) + 0.65 (Stage 2) |
[…]
3.5 Analysis of the computational cost
Generalization is an important ability that allows a DL model to effectively process variant datasets that are not used during training [zhang2021understanding]. To further explore the generalization capability, the Opunake dataset, which is from another 3D survey in New Zealand, is utilized for further TL of T-2.5D and T-3D. Specifically, we extract approximately 1000 and 250 volumes with a shape of 40×40×16 from the full Opunake dataset for training and validation, respectively, and extract a volume with a shape of 224×224×20 as test data. Here, TL for T-2.5D and T-3D is performed in a full fine-tuning manner. The comparisons of the interpolated results before and after the TL are presented in Fig. 14. As indicated by the red boxes in Fig. 14c and 14d, the interpolated results before applying TL contain certain discontinuous events. As illustrated in Fig. 14e and 14f, this issue is alleviated after applying TL. The numerical results have also been evidently improved. Meanwhile, Fig. 15 gives the corresponding residual images and local similarity maps. Compared to Fig. 15a–15d, the lower energy in Fig. 15e–15h indicates better interpolation performance after applying TL.
We plot the 20th line of the test data from Fig. 14 for 2D comparison. As shown in Fig. 16c and 16e, the two interpolated results before applying TL contain discontinuous events as indicated by the red boxes. After applying TL, the continuity and smoothness have been improved, as shown in Fig. 16g and 16i. T-2.5D still achieves comparable performance to that of T-3D, with an even higher SSIM value. In summary, T-2.5D has great potential for effectively interpolating randomly missing traces in 3D seismic data from different survey areas, exhibiting great generalization ability.
4 Discussions
4.1 Testing in the Regularly Missing Case
All the interpolated results in the previous section are dealing with irregularly missing traces. In fact, T-2.5D can also be leveraged to handle regular gaps to increase the acquisition density. To validate this, we further assesse the proposed T-2.5D on the Kerry dataset with 67% regularly missing traces (removing two out of every three traces) in Fig. 17b. As shown in Fig. 17c, 17e, and 17g, the interpolation result of T-2.5D shows superior performance to T-2D. Furthermore, it has achieved even better numerical results than T-3D does under such a high missing rate. The residual image in Fig. 17h further demonstrates that T-2.5D can effectively recover signals with little leakage in the regularly missing scenario.
4.2 Comparison of Trainable Parameters
The number of trainable parameters in a network model is also an important factor for DL-based methods. Models with more trainable parameters can represent more complex functions, but may require more memory and computation [tu2024overview]. We further compare the trainable parameters of the three networks. Notably, the total trainable parameters of T-2.5D are composed of two parts: the ones from the four Transformer encoders in the first stage, and the ones from the remainder of T-2.5D in the second stage. As shown in Table 5, the trainable parameters surge from 2D to 3D. Although T-2D has the fewest trainable parameters, its interpolation performance is also inferior to the other two networks. Although the total trainable parameters of T-2.5D are only 75.34% of those of T-3D, T-2.5D can achieve performance comparable to T-3D, or even surpassing it in certain cases. This finding further supports the lightweight and efficient nature of T-2.5D.
| Method | Trainable parameters |
| T-2D | 212289 |
| T-3D | 1795953 |
| T-2.5D | 34048 (Stage 1) + 1319121 (Stage 2) |
[…]
4.3 Comparisons of inference time
In DL-based seismic data pre-processing, the inference time also plays an important role due to the need for processing massive datasets. Hence, we have recorded the inference time of T-2D, T-3D and T-2.5D in randomly missing cases. As shown in Table 6, the inference time of T-2.5D is even less than that of T-2D. This is because T-2D requires additional processing time for splitting and reassembling the 3D volumes, which is not needed in T-2.5D. Meanwhile, it only requires less than 1/9 the inference time of T-3D. These comparisons demonstrate that T-2.5D is capable of interpolating seismic data rapidly with high quality when faced with large-scale 3D seismic datasets.
| Inference time (s) | |||
| Method | Kerry dataset | Parihaka dataset | Opunake Dataset |
| T-2D | 8.29 | 8.31 | N/A |
| T-3D | 66.23 | 66.12 | 66.04 |
| T-2.5D | 7.14 | 7.04 | 7.07 |
[…]
5 Conclusion
In this paper, we propose a Transformer-based network, named T-2.5D, to accomplish 3D seismic data interpolation tasks. Specifically, a cross-dimensional TL training strategy is utilized to optimize T-2.5D, so as to reduce the computational burdens when interpolating 3D seismic data with Transformer. This strategy consists of two stages. The first stage is the 2D pre-training stage. It generates a pre-trained model by optimizing the Transformer encoders in T-2.5D using a large amount of 2D data patches. The second stage is the 3D fine-tuning stage, in which the 3D SDAs are fine-tuned using a small number of 3D data volumes. In this stage, the SDAs, each placed before a Transformer encoder, can learn abundant spatial correlation information across seismic lines. Extensive experiments demonstrate that the T-2.5D with this cross-dimensional strategy exhibits comparable or even better performance to T-3D, while requiring only small amounts of memory and training time. In summary, the proposed T-2.5D is qualified to replace time-consuming, memory-intensive full 3D training, so as to significantly improve the efficiency of 3D interpolation using Transformer.